I need to prove and identify is the system fulfill the following characteristics (part 1 to part 4)?

yn] =n2x[n +1]

part 1: is it linear?

part 2: is it BIBO stable?

part 3: is it causal?

part 4: is it time invariant?
Answer #1

Ylm) = n² al-n+1) Part (1) all) I Yn) = ² x 1-n+1) 0,,1m) +QnRy in) Ivylm) = 0,max(-+) ta, max (-741) il silm I vm) , in I (1Tym = 5 1K 12 = 0 KE- / unbounded output) not tinit summation hence system is unstable Part 3 Ylm = nhat-n+1) not causal beca